OpenStudy (anonymous):

5x^2+2x-7=0 solve by using the quadratic formula

OpenStudy (diyadiya):

Quadratic formula : \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (diyadiya):

a=5 b=2 c=-7

OpenStudy (anonymous):

i have a math test next tuesday and these are problems i get stuck on

OpenStudy (anonymous):

\[\LARGE x_{1/2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\LARGE ax^2+bx+c=0\]

OpenStudy (anonymous):

can you show me how you plugged those in im lost

OpenStudy (diyadiya):

\[x=\frac{-2 \pm \sqrt{(2)^2-4(5)(-7)}}{2(5)}\]

OpenStudy (anonymous):

what do i do with the 10 at the bottom

OpenStudy (diyadiya):

did you get the numerator?

OpenStudy (anonymous):

ohhhh got it! thanks!

OpenStudy (diyadiya):

I hope you can solve now